Can you help me solve this? x/2 + 1 = x/4
Answer by nyc_function(2741)

(x/2) + 1 = (x/4)
LCD = 4
(x/2) * 4 + 1*4 = (x/4)*4
2x + 4 = x
2x - x = -4
x = -4